OpenStudy (deadshot):

Simplify quantity 12 x minus 96 over quantity x squared minus 4x minus 32 all over quantity 6 x squared minus 6 x minus 12 over 3 x plus 3

OpenStudy (ajprincess):

Is the question this? \[\LARGE\frac{ \frac{ 12x-96 }{ x^2-4x-32 } }{ \frac{ 6x^2-6x-12 }{ 3x+3} }\]

OpenStudy (deadshot):

Yes, it is

OpenStudy (ajprincess):

It is equal to this. \[\LARGE\frac{ 12x-96 }{ x^2-4x-32}*\frac{ 3x+3 }{ 6x^2-6x-12 }\] Nw try to factorise the numerator and denominator .

OpenStudy (deadshot):

So then it's \[\frac{ 36x^2-288 }{ 6x^4-24x^2+384 }\] and thin I factorize it, right?

OpenStudy (anonymous):

6 over open parentheses x plus 4 close parentheses open parentheses x minus 2 close parentheses quantity x plus 4 over 6 open parentheses x minus 2 close parentheses open parentheses x plus 4 close parentheses open parentheses x minus 2 close parentheses over 6 6 open parentheses x plus 4 close parentheses over x minus 2

OpenStudy (anonymous):

those are the anser choices

OpenStudy (anonymous):

answer *